2. A random sample of n=25 measurements is selected from a population with mean μ = 3 and standard deviation = 1. What is the probability that



a. The sample mean is greater than or equal to 3.1?



b. The sample mean is greater to 2.8, but less than or equal to3.2?

1
Expert's answer
2022-05-09T11:09:13-0400

a.


"P(\bar{X}\ge3.1)=1-P(Z<\dfrac{3.1-3}{1/\sqrt{25}})"

"=1-P(Z<0.5)\approx0.3085"

b.


"P(2.8<\bar{X}\le3.2)=P(Z\le\dfrac{3.2-3}{1/\sqrt{25}})"

"-P(Z<\dfrac{2.8-3}{1/\sqrt{25}})=P(Z\le1)-P(Z<-1)"

"\approx0.84134-0.15866\approx0.6827"